OpenStudy (anonymous):

You are given a 2 digit number. 7 times its tens digit plus 6 times its ones digit is equal to 75; 8 times its tens digit minus 1 times its ones digit is equal to 15. What is the number ?

myininaya (myininaya):

sometthing here let x be the 10's digit let y be the 1's digit 7x+6y=75 what's the other equation?

OpenStudy (anonymous):

wouldnt it be : 8x - 1 = 15? Wait that might be totally wrong.. :P

OpenStudy (anonymous):

ok but does anyone know the answer?

myininaya (myininaya):

i did the last problem for you. shouldn't you learn how to do then for me to give you the answers

OpenStudy (anonymous):

i just need one more example

myininaya (myininaya):

close I<fun

myininaya (myininaya):

fun is close help her with the equation should put down

OpenStudy (anonymous):

oh ^^ thx i figured it wasnt totally rite :P lol and call me dolly!!

myininaya (myininaya):

shes missing one little thing

OpenStudy (anonymous):

?

myininaya (myininaya):

what does y represent?

OpenStudy (anonymous):

the answer

myininaya (myininaya):

above i said let y represent the 1's digit remember?

OpenStudy (anonymous):

ohh ok

myininaya (myininaya):

8 times its ten's digit minus 1 times its one's digit is equal to 15 its one's digit in symbols is... 8 * x - 1 * y = 15

myininaya (myininaya):

8x-y=15

OpenStudy (anonymous):

so the answer is y= 8x+25

OpenStudy (anonymous):

so what is the answer?

myininaya (myininaya):

we haven't got there yet we have two equations 8x-y=15 and 7x+6y=75

OpenStudy (anonymous):

ok

myininaya (myininaya):

try to solve that first one for either x or y and let me know what you come up with

OpenStudy (anonymous):

y=8x-15

myininaya (myininaya):

good now plug that into the y in the other equation so you would 7x+6(_________)=75 what goes in that blank?

OpenStudy (anonymous):

7x+42-90=75

myininaya (myininaya):

7x+6(8x-15)=75 7x+48x-80=75 now see if you can solve this for x

OpenStudy (anonymous):

55x=155

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

you made a mistake 6 times 15 is 90

OpenStudy (anonymous):

is that the correct answer?

myininaya (myininaya):

ok then it would be 55x=165 which would make x=3

myininaya (myininaya):

but thats the 10's digit whats the one's digit

OpenStudy (anonymous):

its 39

myininaya (myininaya):

so if x=3, then y=8(3)-15=24-15=9 so yes the answer is 39

OpenStudy (radar):

myininaya just received a medal from radar

OpenStudy (anonymous):

radar just recieved a medal from dolly!!